What is the volume and surface area of a cube with an edge length of 4 cm?
1
Expert's answer
2011-07-27T06:26:22-0400
The volume of a cube with the edge a = 4 cm is
V = a3 = 43 = 64 cm3.
The surface area consists of 6 squares with the side of 4 cm, hence
S = 6* a2 = 6*42 = 96 cm2.